The mean life of a belted tire is given as 45000 km. The standard deviation is 2250 km.
What percent of the tires will have a life that exceeds 47250 km?
What percent of the tires will last longer than 40500 km?

Answer by stanbon(75887)

The mean life of a belted tire is given as 45000 km. The standard deviation is 2250 km.
What percent of the tires will have a life that exceeds 47250 km?
What percent of the tires will last longer than 40500 km?
--------------------------
Find the z-score for 47250 and for 40500.
z(47250) = (47250-45000)/2250 = 1
P(z> 1 ) = 0.15866...
-----------------------------
z(40500) = (40500-45000)/2250 = -4500/2250 = -2
P(z>-2) = 0.9772499...
======================
Cheers,
Stan H.